# From Social Division to Cohesion with AI Message Suggestions in Online Chat Groups

### v1

@@ Abstract
Social cohesion is difficult to sustain in societies marked by opinion diversity, particularly in online communication. As large language model (LLM)-driven messaging assistance becomes increasingly embedded in these contexts, it raises critical questions about its societal impact. We present an online experiment with 557 participants who engaged in multi-round discussions on politically controversial topics while freely reconfiguring their discussion groups. In some conditions, participants received real-time message suggestions generated by an LLM, either personalized to the individual or adapted to their group context. We find that subtle shifts in linguistic style during communication, mediated by AI assistance, can scale up to reshape collective structures. While individual-focused assistance leads users to segregate into like-minded groups, relational assistance that incorporates group members’ stances enhances cohesion through more receptive exchanges. These findings demonstrate that AI-mediated communication can support social cohesion in diverse groups, but outcomes critically depend on how personalization is designed.
@@
1 Introduction
Social cohesion remains a persistent challenge in societies characterized by opinion diversity [1, 2, 3]. Conversations across differences are fraught: people struggle to engage constructively with those who hold opposing views, and stereotypes bias perceptions of outgroup members [4, 5]. The challenges individuals face in conversation are reinforced by structural dynamics, as homophily drives clustering with similar others and limits cross-cutting ties at scale [6, 7, 8]. Online platforms intensify this micro–macro loop by lowering inhibitions in communication, accelerating interpersonal clustering, and amplifying these effects at scale [9, 10, 11, 12, 13]. The consequences—from political polarization to breakdowns in global coordination—underscore the urgency of understanding how opinion diversity is managed within online environments [14, 15, 16, 17, 18].
Recent advances in artificial intelligence (AI) introduce a new layer of influence to this challenge [19, 20]. Large language models (LLMs) are now widely integrated into everyday online communication through messaging assistance tools [21, 22, 23]. While such assistance can reduce the burden of writing [24], conventional personalization features risk amplifying stance differences among people with diverse views [25, 26]. Conversely, some designs for AI-mediated communication have been proposed to foster interpersonal trust and build consensus across lines of disagreement [27, 1, 22, 28, 29]. Yet, it remains unclear how these micro-level interventions can cascade upward to shape macro-level changes in human group composition. How can AI message suggestions foster social cohesion rather than exacerbate division in diverse populations?
@@ 1 Introduction
Here, we address this question through a preregistered experiment conducted on a customized online chat platform (see Methods for details) [30]. Participants ($N=557$), recruited from Prolific [31], were randomly assigned to sessions of 6 to 15 people ($M=9.42$, $s.d.=2.57$), subdivided into groups of 2 to 4 (Fig. 1). Each group discussed a politically controversial topic selected to maximize opinion diversity for that session, based on participants’ pre-task survey responses (see Extended Data Table 1 for survey items) [27]. Participants were able to update their survey stance during the session, though this occurred rarely ($6.40\%$ of cases).
@@
Participants were instructed to engage in group conversations, but were not asked to reach consensus, reflecting the open-ended and non-teleological nature of informal online discussions [32, 14]. After each three-minute discussion, participants decided whether to remain in their current group, join another, or create a new one. If only one participant remained in a group, they continued the discussion with a chatbot to ensure continuous engagement. This procedure was repeated for four rounds, followed by a final switching decision that determined the Round 5 group composition used for structural analysis.
@@
Within this setup, we manipulated the presence and type of AI assistance in message writing through three experimental conditions [27]. In the no-assistance condition, participants composed all messages themselves and did not receive any AI-generated suggestions. In the other two conditions, participants received real-time message suggestions generated by OpenAI’s GPT-4o model as the group chat unfolded. Although suggestions were tailored to each user based on their survey responses and subsequent updates in both conditions, the personalization context differed between them (see Methods; full prompts appear in SI.).
@@
In the individual assistance condition, the model was prompted to help participants articulate their stance on the conversation topic, aligning suggestions with each individual’s viewpoint. In contrast, in the relational assistance condition, the modelincorporated not only the participant’s stance but also those of their group members, with the goal of fostering mutual understanding. As a result, individual-focused suggestions primarily supported self-expression, whereas relationally aware suggestions more often conveyed receptive language and positive sentiment, even in the presence of disagreement (Extended Data Table 2), as confirmed by quantitative analyses (Extended Data Fig. 1). These suggestions could be sent with a single click, edited, or ignored in favor of self-written messages.
@@
Following the preregistered protocol [30], we excluded sessions with fewer than six participants or with more than 25% dropout. The final dataset includes 20 sessions per condition (60 sessions total), comprising 557 participants overall (187 in the no assistance condition, 182 in the individual assistance condition, and 188 in the relational assistance condition). Each participant completed only one session. Per-session participant counts, dropout rates, and topic assignments did not differ significantly across conditions (Chi-square tests for participant counts: $p=0.946$; dropout rates: $p=0.538$; Fisher’s exact test for topic distributions: $p=0.533$).
@@
We evaluate the effects of AI assistance on group composition and social cohesion, measured through stance assortativity—a session-level indicator of how strongly group composition reflects similarity in participants’ opinions. Individuals are treated as connected if they belong to the same group [33], and assortativity is computed from their stances on the discussion topic (3-3 = totally disagree to +3+3 = totally agree) [34]. Higher assortativity values indicate stronger clustering of like-minded individuals (i.e., greater social division and lower cohesion), whereas lower values indicate more cross-cutting ties (i.e., lower division and greater cohesion) (Fig. 1). Because random assignment produced variation in initial group composition, we analyze changes relative to the first-round baseline. To examine the mechanisms underlying these structural shifts, we also analyze conversational dynamics, including message volume and sentiment, with the latter measured using VADER sentiment analysis [35].
Figure 1: Illustration of group-composition dynamics. Each session begins with 2 to 4 groups, each consisting of 2 to 4 members, shown as gray clusters. Stance assortativity is calculated at the session level from participants’ stances on the discussion topic, represented by node colors. As participants switch groups across rounds, group composition evolves, either increasing social division (positive assortativity) or fostering cohesion (negative assortativity).
2 Results
2.1 Effects of AI Assistance on Social Cohesion
@@
We found that stance assortativity remained stable or increased slightly across rounds in the no-assistance condition, but decreased substantially in the relational assistance condition (Fig. 2). After four rounds, assortativity decreased by an average of 0.135 points on the –1 to 1 scale under relational assistance. This reduction was statistically significant relative to the no-assistance condition ($p=0.004$; regression analysis; Extended Data Table 3). In contrast, changes in assortativity under individual assistance did not differ significantly from the control ($p=0.608$). Relational assistance promoted more cross-cutting group composition, while individual assistance did not alter baseline patterns of clustering.
To further understand the effects of AI assistance on structural dynamics, we examined the number of conversation partners, the number of conversation groups, within-group stance distance, and between-group stance distance across interaction rounds. Participants, on average, increased their conversation partners without AI assistance ($p<0.001$ in the no-assistance condition; Fig. 3a). This trend was observed across all conditions but it was amplified under individual assistance: the increase in conversation partners was significantly larger than in the control ($p=0.028$; Extended Data Table 3). Relational assistance did not differ significantly from the control ($p=0.712$).
The number of groups remained stable across rounds in the control condition ($p=0.280$), and neither AI condition altered this pattern (individual: $p=0.317$; relational: $p=0.142$; Fig. 3b). Combined with the observed increase in individual connectivity (Fig. 3a), this pattern indicates that participants tended to consolidate into a large group, while some drifted away from the largest component. This consolidation was most pronounced in the individual assistance condition, which exhibited the highest overall connectivity.
Stance-based group composition, measured by stance assortativity (Fig. 2), can also be characterized by average stance differences within and between groups. We found that within-group stance distance remained stable or decreased slightly in the no-assistance and individual assistance conditions. In the relational assistance condition, within-group stance distance marginally increased relative to the no-assistance condition ($p=0.070$; Fig. 3c). In contrast, between-group stance distance increased significantly over rounds in the control condition ($p=0.017$), and this separation was further amplified under individual assistance ($p=0.004$; Fig. 3d). Relational assistance, however, did not differ significantly from the control in between-group stance distance ($p=0.350$). In summary, relational assistance increased ideological diversity within groups, whereas individual assistance widened separation between groups.
@@
2.2 From AI Assistance to Conversational Receptiveness
AI assistance in this study operated solely at the level of individual message writing, without directly shaping or suggesting group formation. This implies that a mediating process likely links individual use of AI assistance to macro-level changes in social cohesion. We therefore examined participants’ AI suggestion use and group conversational dynamics as potential mediators.
In the AI-assisted conditions, participants continuously received message suggestions as the group chat unfolded and occasionally chose to send them. Overall, $45.1\%$ of messages in the individual condition and $37.0\%$ in the relational condition originated from GPT-4o. Most of these AI messages were sent without edits ($96.3\%$ of suggestions used). This AI use increased participants’ overall communication volume (Fig. 4a). When suggestions were available, participants sent more messages, an effect especially pronounced in the individual condition (up 37.9% in the individual condition; up $20.0\%$ in the relational condition; $p\leq 0.001$ for both). They, however, produced fewer self-written messages than in the no-assistance condition: the number of self-written messages decreased by $24.2\%$ in the individual assistance condition and by $24.5\%$ in the relational assistance condition ($p<0.001$ for both).
@@ 2.2 From AI Assistance to Conversational Receptiveness
AI assistance also altered participants’ communication style (Fig. 4b). Relational assistance, which incorporated social awareness by design, generated more emotionally positive and receptive messages than individual assistance or self-written messages (Extended Data Fig. 1). As participants occasionally used these suggestions, the overall emotional valence of their messages increased; VADER scores, which quantify the sentiment polarity of text—with higher scores reflecting more positive emotional tone [35], rose from $0.113$ without assistance to $0.249$ with relational assistance ($p<0.001$). Notably, even participants’ self-written messages in the relational assistance condition were significantly more positive than those in the no-assistance condition ($p=0.006$). We confirmed that this increase in emotional positivity was correlated with other dimensions of conversational receptiveness [36, 37], including greater politeness (Extended Data Fig. 2) and lower toxicity (Extended Data Table 4).
Individual-level changes in messaging with AI assistance shaped group-level conversational dynamics (Extended Data Table 5). In the absence of AI assistance, groups had less communication as rounds progressed ($p=0.048$; Fig. 4c). When individual message suggestions were available, however, groups maintained communication volume, significantly reversing the decline ($p=0.008$). On the other hand, individual assistance did not affect group-level emotional valence ($p=0.206$; Fig. 4d). Relational assistance exhibited the opposite pattern: the volume of the message did not differ from the default ($p=0.664$), but baseline emotional valence was significantly higher ($p=0.004$) and remained stable across rounds.
@@
Regarding the individual–group effect, it is important to note that contributions to group discussions were not evenly distributed across participants. Unlike dyadic conversations driven by reciprocal turn-taking, group discussions are often shaped by a few members who contribute more frequently and, in doing so, set the conversational tone [38, 39]. We find that this heterogeneity was strongly associated with AI usage: participants who relied more heavily on AI message suggestions sent more messages to their groups. Figure 5 shows a strong positive correlation between AI suggestion use and message volume in the individual assistance condition (Spearman’s rank correlation $r=0.536$; $p<0.001$) and a moderate positive correlation in the relational assistance condition ($r=0.387$; $p<0.001$). This pattern likely reflects the reduced effort required to send AI-generated messages, enabling heavier users to contribute at a higher rate. As a result, the linguistic tone embedded in AI-generated suggestions (Extended Data Fig. 1) disproportionately influenced group conversations through these more AI-dependent participants.
2.3 From Conversational Receptiveness to Social Cohesion
@@
After each discussion round, on average $32.6\%$ of participants left their groups to join another or create a new one. There were no significant differences in group-switching rates across conditions ($31.5\%$ in the control; $33.3\%$ in the individual condition; $33.1\%$ in the relational condition; one-way ANOVA, $p=0.553$). This suggests that the system-level differences observed in Figs. 2 and 3 reflect not the overall likelihood of switching groups, but by which groups they chose to leave and join.
Participants’ experiences during group discussions likely influenced whether they stayed in their current group or chose a new one, thereby linking micro-level interaction patterns to macro-level group composition Fig. 2. To examine this, we analyzed whether each participant’s message counts and sentiment predicted the stance distance between that participant and their group members in the next round, while controlling for their current stance distance (Fig. 6a). Linear mixed models were used for this analysis, with effects estimated separately for participants who stayed in the same group and those who switched groups (see Methods for details).
@@
Figure 6: Effects of message sentiment and volume on ego-centric stance distance. (a) Example of changes in ego-centric stance distance across rounds. One member left and two joined (indicated by bold lines) from Round $t$ to Round $t+1$. Node color indicates each member’s stance (see Fig. 1). (b) Heatmap shows predicted between-round change in ego-centric stance distance for participants who remained in the same group, estimated from a linear mixed model (see Extended Data Table 6). The contour line marks the transition point ($\Delta=0$) between reduced and increased diversity. Axes represent average sentiment (VADER score) and number of messages sent. Subfigures show the density distributions of message sentiment and counts across AI assistance conditions.
Figure 6b presents the estimated results for participants who remained in the same group (see Extended Data Table 6 for model estimates). We found that emotional valence increased each participant’s stance distance when interacting with higher message counts ($p=0.031$). Given that participants rarely updated their stance during the conversation (6.40% of cases), this increase likely reflects the formation of more heterogeneous connections within the group rather than the emergence of polarization (Fig. 6a). In other words, when individuals sent more emotionally positive messages, their groups tended to attract a more ideologically diverse set of members. By contrast, for participants who switched groups, we observe no consistent relationship between prior communication and the composition of their subsequent group (Extended Data Table 6).
@@
Finally, we examined the entire causal pathways using structural equation modeling [40] at the session level (Extended Data Table 7). The path diagram in Figure 7 reveals distinct mechanisms for individual and relational assistance. Individual assistance increased message volume, whereas relational assistance strongly increased message sentiment. Higher message volume alone was associated with greater stance assortativity, but its interaction with positive sentiment predicted lower assortativity in the subsequent round, reflecting the emergence of more heterogeneous group configurations. The effects of assistance on assortativity were fully mediated by conversational dynamics, indicating that structural changes emerge indirectly through interaction patterns rather than direct intervention.

### v2

# AI-Mediated Communication Reshapes Social Structure in Opinion-Diverse Groups

@@ Abstract
Group segregation or cohesion can emerge from micro-level communication, and AI-assisted messaging may shape this process. Here, we report a preregistered online experiment (N = 557 across 60 sessions) in which participants discussed controversial political topics over multiple rounds and could freely change groups. Some participants received real-time message suggestions from a large language model (LLM), either personalized to their stance (“individual assistance”) or incorporating their group members’ perspectives (“relational assistance”). We find that small variations in AI-mediated communication cascade into macro-level differences in group composition. Participants with individual assistance send more messages and show greater stance-based clustering, whereas those with relational assistance use more receptive language and form more heterogeneous ties. Hybrid expressive processes—jointly produced by humans and AI—can reshape collective organization. The patterns of structural division and cohesion depend on how AI incorporates users’ interaction context.
@@
Understanding how micro-level communication patterns accumulate into macro-level group segregation or cohesion is a central question in social and behavioral science [1, 2, 3]. Conversations across differences are often asymmetric: people find it difficult to engage constructively with those who hold opposing views [4, 5], and stereotypes bias perceptions of outgroup members [6]. These interpersonal tendencies interact with structural forces—particularly homophily and clustering—to produce emergent segregation, even without explicit hostility [2, 7, 8]. Online platforms can intensify these dynamics through lowered inhibitions [9], emotion-amplified diffusion [10], and algorithmic or behavioral clustering processes [11, 12, 13]. While the forces that produce social division are well theorized and empirically documented, far less is known about the micro-level conversational mechanisms that can instead generate cohesion in ideollogically diverse groups [14, 15, 16].
Recent advances in artificial intelligence (AI) introduce a new layer of influence to these micro-macro dynamics [17]. Large language models (LLMs) are increasingly integrated into everyday communication through messaging assistance tools [18, 19]. These tools effectively create hybrid human–AI expressive processes, in which algorithmic suggestions subtly shape how people articulate their views [20, 21, 22, 23]. While such assistance can reduce the burden of writing [24], conventional personalization features may unintentionally reinforce users’ ideological viewpoints and deepen intergroup divisions [25, 26]. In contrast, recent work shows that assistance incorporating the perspectives of interaction partners can improve cross-group receptiveness [27, 28, 29, 30]. Yet it remains unclear whether—and how—such micro-level shifts in AI-mediated expression scale up to shape macro-level patterns of group composition in hybrid communication environments.
@@
To examine these dynamics, we conducted a preregistered experiment on an online chat platform (see Methods for details) [31]. Participants ($N=557$), recruited from Prolific [32], were randomly assigned to sessions of 6 to 15 people ($M=9.42$, $s.d.=2.57$), subdivided into groups of 2 to 4 (Fig. 1). Groups discussed a politically controversial topic selected to maximize opinion diversity within each session, based on pre-task survey responses (Extended Data Table 1) [27]. Participants could update their stance during the session, though this occurred rarely ($6.40\%$ of cases).
@@
Participants were instructed to engage in open-ended discussions, without any requirement to reach consensus, mirroring the non-teleological nature of informal online conversations [33, 13]. After each three-minute discussion, participants chose whether to remain in their current group, join another, or create a new one. If a group had only one participant, they continued the discussion with a chatbot to ensure continuous engagement. This process was repeated for four rounds, followed by a final switching decision that determined the final group configuration used for structural analysis.
@@
Within this setup, we manipulated the presence and construction of AI assistance through three conditions [27]. In the no-assistance condition, participants composed all messages themselves. In the other two conditions, participants received real-time message suggestions generated by OpenAI’s GPT-4o model, which they could send directly with one click, edit, or ignore. Suggestions were personalized to each participant based on their survey-reported stance and the unfolding conversation, but the nature of this personalization differed across the two assisted conditions.
@@
In the individual assistance condition, the model was prompted to help participants articulate their own stance on the discussion topic, supporting self-expression. In the relational assistance condition, the model was prompted to incorporate both the participant’s stance and those of their group members, with the emphasis on fostering mutual understanding. Individual and relational prompts were matched in length and complexity and differed only in instructional framing (see Methods; full prompts appear in SI). As a consequence, individual-focused suggestions emphasized self-expression, whereas relationally framed suggestions adapted to group composition and conveyed more positive sentiment, even in the presence of disagreement (Extended Data Table 2), a pattern confirmed by quantitative analyses (Extended Data Fig. 1). This variation enabled a causal test of how distinct expressive framings influence group-level conversation and structural dynamics.
@@
Following the preregistered protocol [31], we excluded sessions with fewer than six participants or with more than 25% dropout. The final dataset includes 20 sessions per condition (60 sessions total), comprising 557 unique participants overall (187 in the no assistance condition, 182 in the individual assistance condition, and 188 in the relational assistance condition). Per-session participant counts, dropout rates, and topic assignments did not differ significantly across conditions (participant counts: $p=0.946$; dropout: $p=0.538$; topics: $p=0.533$).
@@
We evaluate effects on group composition using stance assortativity, a structural measure of how strongly group membership reflects similarity in opinions (Fig. 1). Individuals are defined as connected if they belong to the same group [34], and assortativity is computed from their stance values (3-3 = totally disagree to +3+3 = totally agree) [35]. Higher assortativity indicates stronger clustering of like-minded individuals, whereas lower assortativity indicates more cross-cutting ties and greater cohesion. Because random assignment produced variation in initial group composition, we analyze changes relative to the first-round baseline. To probe the mechanisms underlying these structural shifts, we also examine conversational dynamics, including message volume and sentiment (measured using VADER [36]).
Figure 1: AI-mediated communication and group reconfiguration. This schematic illustrates the experimental setup and the conceptual process by which group composition can reorganize during the study. Nodes represent participants (with colors indicating their stance), and edges denote co-membership in discussion groups. In the AI-assisted conditions, message suggestions generated by a LLM are tailored to each participant’s stance and conversational context, creating hybrid human–AI expressive processes. Stance assortativity is computed at the session level and reflects how strongly group membership clusters like-minded individuals. As participants switch groups, group composition can reorganize in ways that increase or decrease assortativity, corresponding to greater social division or cohesion.
Results
Effects of AI Assistance on Group Reconfiguration
@@ Results
We found that stance assortativity remained stable or increased slightly across rounds in the no-assistance condition, but decreased substantially under relational assistance (Fig. 2). After four rounds, assortativity decreased by an average of 0.135 points (on the –1 to 1 scale) in the relational assistance condition, a statistically significant reduction relative to the no-assistance control ($p=0.004$; regression analysis; Extended Data Table 3). In contrast, changes in assortativity under individual assistance did not differ significantly from the control ($p=0.608$). Thus, relational assistance promoted more heterogeneous, cross-cutting ties, while individual assistance did not alter baseline assortativity.
To further characterize how AI assistance shaped structural dynamics, we examined changes in the number of conversation partners and the number of conversation groups across rounds. Participants, on average, increased their number of conversation partners even without assistance ($p<0.001$ in the no-assistance condition; Fig. 3a). This increase was significantly larger under individual assistance than in the non-assistance control ($p=0.028$; Extended Data Table 3), whereas relational assistance did not differ from the control ($p=0.712$). The number of groups remained stable across rounds in the control condition ($p=0.280$), and neither form of assistance significantly altered this pattern (individual: $p=0.317$; relational: $p=0.142$; Fig. 3b). Given that individual connectivity increased while the number of groups remained constant, participants tended to consolidate into a single large group while some drifted away from the largest component. This consolidation was most pronounced under individual assistance, which exhibited the highest overall connectivity.
We further examined structural diversity by complementing stance assortativity with group-based measures of average stance differences within and between groups. Within-group stance distance remained stable or decreased slightly in the no-assistance and individual assistance conditions. Under relational assistance, within-group stance distance marginally increased relative to the no-assistance condition ($p=0.070$; Fig. 3c). Between-group stance distance increased significantly over rounds in the control condition ($p=0.017$), and this separation was further amplified under individual assistance ($p=0.004$; Fig. 3d). Relational assistance did not differ significantly from the control in between-group stance distance ($p=0.350$). Taken together, individual assistance produced groups that became more internally uniform and more externally separated, whereas relational assistance did not increase between-group separation and instead diversified group memberships.
@@
From AI Assistance to Conversational Dynamics
AI assistance in this study operated only at the level of individual message composition and never directly involved group reconfiguration. Accordingly, any macro-level structural differences must emerge through processes unfolding within the conversations themselves. To identify these processes, we examined participants’ use of AI suggestions and the resulting conversational dynamics as potential mediators linking expressive framing to group-level outcomes.
In the AI-assisted conditions, participants continuously received message suggestions as the conversation unfolded and occasionally chose to send them. Overall, $45.1\%$ of messages in the individual condition and $37.0\%$ in the relational condition originated from GPT-4o. Moreover, 96.3% of these suggestions were sent without modification, indicating a high degree of reliance on and integration of AI-generated language within the hybrid human–AI expressive process.
When suggestions were available, participants sent more total messages (Fig. 4a)—an effect especially pronounced in the individual condition (up to a 37.9% increase in total messages; $p<0.001$) and to a lesser extent in the relational condition (up to a 20.0% increase; $p=0.001$). On the other hand, participants produced fewer self-written messages under AI assistance. Self-written message volume decreased by 24.2% in the individual condition and by 24.5% in the relational condition ($p<0.001$ for both).
@@
AI assistance also altered participants’ communication style (Fig. 4b). Relational assistance, which incorporated social awareness by design, generated more emotionally positive and receptive suggestions than individual assistance or self-written messages (Extended Data Fig. 1). Participants who received these suggestions sent messages with significantly more positive emotional tone. VADER scores, which quantify the sentiment polarity of text, rose from $0.113$ to $0.249$ under relational assistance ($p<0.001$) (higher scores reflecting more positive emotional tone [36]). Notably, even participants’ self-written messages in the relational assistance condition were significantly more positive than those in the no-assistance condition ($p=0.006$). We confirmed that this increase in emotional positivity was correlated with other dimensions of conversational receptiveness [37, 38], including greater politeness and lower toxicity (Extended Data Fig. 2).
Individual-level changes in messaging with AI assistance scaled up to shape group-level conversational dynamics (see Extended Data Table 4 for illustrative discussion excerpts). In the absence of AI assistance, groups communicated less as rounds progressed ($p=0.048$; Fig. 4c). Under individual assistance, groups maintained conversation volume and significantly reversed this downward trend ($p=0.008$), while conversational sentiment remained unchanged ($p=0.206$). Relational assistance, in contrast, did not alter conversation volume trajectories ($p=0.664$), but produced significantly more positive baseline sentiment ($p=0.004$), which remained stable across rounds (Fig. 4d; see Extended Data Table 5 for full statistical results).
@@
Regarding the individual–group link, contributions to group discussions were not evenly distributed across participants. Unlike dyadic conversations driven by reciprocal turn-taking, group discussions are often dominated by members who contribute more frequently and, in doing so, set the conversational tone and norms [39, 40]. Figure 5 shows that this heterogeneity was strongly associated with AI usage: participants who relied more heavily on AI suggestions sent more messages to their groups (Spearman’s $r=0.536$ under individual assistance; $r=0.387$ under relational assistance; $p<0.001$ for both). This pattern likely reflects that the reduced effort required to send AI-generated messages enables heavier users to contribute at a higher rate. As a consequence, uneven reliance on AI created asymmetric influence within group discourse, amplifying the expressive impact of the most AI-dependent participants.
From Conversational Dynamics to Group Reconfiguration
@@
After each discussion round, an average of $32.6\%$ of participants left their groups to join another group or create a new one. Group-switching rates did not differ significantly across conditions ($31.5\%$ in the no-assistance condtion; $33.3\%$ in the individual assistance condition; $33.1\%$ in the relational assistance condition; one-way ANOVA, $p=0.553$). This indicates that the system-level differences observed in Figs. 2 and 3 were not driven by whether participants switched groups, but rather by where they moved. And participants’ conversational experiences likely influenced their subsequent group-switching decisions.
To examine this linkage, we analyzed whether a participant’s message volume and sentiment in a given round predicted changes in their stance distance from their group members in the next round, controlling for their current stance distance (Fig. 6a). Figure 6b shows that more positive message sentiment increased next-round stance distance when paired with higher message volume ($p=0.031$; see Extended Data Table 6 for full model results). Because participants rarely updated their own stance relative to the pre-task survey (6.40% of cases), these changes primarily reflect shifts in group composition rather than ideological movement by individuals. In effect, participants who expressed more positive sentiment tended to end up in groups that became more ideologically diverse in the subsequent round. This provides a micro-level mechanism consistent with the reduction in stance assortativity observed under relational assistance (Fig. 2).
@@
Figure 6: Effects of message sentiment and volume on ego-centric stance distance. (a) Example of changes in ego-centric stance distance across rounds. From Round $t$ to Round t+1t{+}1, one member left and two members joined the group (bold edges). Node color indicates each member’s stance (see Fig. 1). (b) Predicted between-round change in ego-centric stance distance for participants who remained in the same group, estimated using a linear mixed model (Extended Data Table 6). The contour line marks the transition point ($\Delta=0$) between reduced and increased stance distance. Axes represent average sentiment (VADER score) and number of messages sent. Marginal density plots show the observed distributions of message sentiment and message counts across the three AI assistance conditions.
@@
Finally, we examined the full causal pathways using structural equation modeling [41] at the session level (Extended Data Table 7). The path diagram in Figure 7 reveals distinct mechanisms for individual and relational assistance. Individual assistance mainly increased message volume, whereas relational assistance increased message sentiment. Higher message volume alone was associated with greater stance assortativity, but its interaction with positive sentiment predicted lower assortativity in the subsequent round, reflecting the emergence of more heterogeneous group configurations. The effects of assistance on assortativity were fully mediated by conversational dynamics (i.e., no significant direct effects), indicating that structural changes arose indirectly through interaction patterns rather than direct intervention. Consistent with our theoretical framing, neither form of assistance directly altered assortativity; instead, structural reorganization resulted from how hybrid expressive processes reshaped local conversational dynamics.